OpenStudy (anonymous):

First to solve= free medal. Derive the continuous compounding interest formula (A=Pert ) from the compound interestert formula A=P(1 + r/n)nt

OpenStudy (zarkon):

\[(1+r/n)^n=e^{n\ln(1+r/n)}\] now take the limit as \[n\to \infty\]